The length of a rectangle is 3 feet less than twice the width of the rectangle. If the perimeter of the rectangle is 414 feet, find the width and the length.
Answer by kainharen(2)

Perimeter of rectangle = 2(l+b)
so according to question length of rectangle is 3 feet less than twice the width of the rectangle
so
let the width of rectangle=x
then the length of rectangle= 2x-3
2[(2x-3)+x]=414
4x-6+2x=414
6x=414+6
x=420/6
x=70
then,
length of rectangle= 2x-3=140-3=137 feets.
Widt of rectangle= 70 feets.